# MLAAN: Scaling Supervised Local Learning with Multilaminar Leap Augmented Auxiliary Network

### v4

@@ Introduction
End-to-end (E2E) backpropagation, a mainstream method for training deep neural networks, has achieved remarkable success in image segmentation (Wang, Peng, and Liu 2018), target tracking (Bao et al. 2022), visual sound separation (He et al. 2021), and other fields. Despite these accomplishments, E2E learning has its drawbacks. It necessitates retaining the complete computational graph and all intermediate activations post each local module’s computation (Wang et al. 2021a), increasing GPU memory usage. As network depth and model complexity grow, this memory demand can severely constrain the training on high-resolution data inputs (Wu et al. 2021). Additionally, the E2E structure delays parameter updates in initial layers until the entire forward and backward propagation cycles conclude, causing inefficiency (Jaderberg et al. 2017a), (Salimans et al. 2017). This sequential gradient processing also impedes model parallelization (Löwe, O’Connor, and Veeling 2019a), (Belilovsky, Eickenberg, and Oyallon 2020) and raises biological plausibility concerns due to its reliance on cross-layer backpropagation (Crick 1989), (Shen, Zhao, and Zeng 2022).
Due to the inherent flaws of E2E training (Illing et al. 2021a), (Xiong, Ren, and Urtasun 2020a), some researchers are exploring local learning as an alternative. Local learning divides the network into gradient-isolated local modules and designs auxiliary networks to assist their training (Bengio et al. 2006). This method eliminates the need for sharing gradients and activation parameters between modules, reducing GPU memory requirements (Wang et al. 2021a). In addition, local learning uses local error signals, which align more closely with biological feasibility principles. However, the existing focus on local features and the lack of information exchange between local modules significantly lower overall performance than that of E2E training (Mostafa, Ramesh, and Cauwenberghs 2018). Despite some recent improvements in the performance of local learning (Belilovsky, Eickenberg, and Oyallon 2020), (Nøkland and Eidnes 2019), current methods still have myopia issues and insufficient global feature representation, limiting their application in precision-critical tasks. Therefore, the crucial goal of local learning is to ensure that each local module receives ample global information and facilitates information exchange between modules. Maintaining GPU memory efficiency while improving global performance is essential for making local learning a viable alternative to E2E training.
@@
Local learning is considered a promising alternative to traditional E2E methods (Nøkland and Eidnes 2019). However, the lack of local inter-module interaction and poor global performance has hindered its development (Bengio et al. 2006). To improve global performance, some scholars have used self-supervised loss to preserve features during local module propagation (Illing et al. 2021b). For example, Y. Xiong et al. (Xiong, Ren, and Urtasun 2020b) designed SimCLR loss for Local Contrastive Learning. Sindy et al. (Löwe, O’Connor, and Veeling 2019b) applied Comparative Predictive Coding loss to reduce feature loss in Greedy InfoMax. These methods perform well with a few network blocks but struggle with many blocks. Adrien et al.’s method (Journé et al. 2022), combining Hebbian learning with local learning, addresses this issue but faces challenges on large datasets like ImageNet (Deng et al. 2009). Methods like InfoPro (Wang et al. 2021a) and DGL (Belilovsky, Eickenberg, and Oyallon 2020) have improved local learning performance on large datasets such as ImageNet (Deng et al. 2009). However, due to gradient isolation, their performance still falls short of E2E methods. Enhancing inter-module communication to match or surpass E2E performance remains a critical issue.
@@
BP(ResNet-32)=19.35, BP(ResNet-110)=19.67
SVHN DGL 4.83 5.05 5.12 5.36
@@
Results on Image Classification Benchmarks. We evaluate MLAAN’s performance on CIFAR-10 (Krizhevsky, Hinton et al. 2009), SVHN (Netzer et al. 2011), and STL-10 (Coates, Ng, and Lee 2011) datasets, with results in Table 1.
On CIFAR-10 (Krizhevsky, Hinton et al. 2009), increasing local modules enhances MLAAN’s benefits. For ResNet-32 (k=16) (He et al. 2016), the Test Error decreases from 14.08 and 12.93 to 6.36 and 6.37, marking a 55% and 51% improvement over the BP (Rumelhart et al. 1985) baseline. For ResNet-110 (k=55) (He et al. 2016), MLAAN with DGL (Belilovsky, Eickenberg, and Oyallon 2019) and InfoPro (Wang et al. 2021b) improves performance by 61% and 56%, respectively. On STL-10 (Coates, Ng, and Lee 2011), MLAAN achieves over 30% and 34% improvements under the same situations, while on SVHN (Netzer et al. 2011), gains exceed 42% and 49%.
@@
These results demonstrate MLAAN’s ability to improve deep learning models, bridge performance gaps, and outperform BP even with more local modules, reinforcing its effectiveness in improving accuracy and robustness in complex and deep architectures.
Memory Consumption. We investigate the GPU memory usage of BP (Rumelhart et al. 1985) and MLAAN on CIFAR-10 (Krizhevsky, Hinton et al. 2009) and ImageNet (Deng et al. 2009), with results in Table 3.
For ResNet-32 (k=16) (He et al. 2016), usage decreases by 22.3% compared to BP (Rumelhart et al. 1985). For ResNet-110 (k=55) (He et al. 2016), the reduction is 70.2%. ResNet-34, ResNet-101, and ResNet-152 (He et al. 2016) also show reductions of 15.4%, 13.1%, and 12.5%, respectively, compared to BP (Rumelhart et al. 1985). These findings demonstrate that MLAAN performs better with less GPU memory usage than BP (Rumelhart et al. 1985).
@@
Comparison of Different Module Combinations. We conduct experiments with ResNet-32 (K=16) (He et al. 2016) as the backbone, using DGL (Belilovsky, Eickenberg, and Oyallon 2019) as the baseline, to evaluate MLM, LAM, and MLAAN. The results are summarized in Table 4.
@@
Comparative Analysis of Unfair Epochs. Undeniably, the utilization of MLAAN introduces additional Wall Time, prompting us to conduct a comparative analysis of unfair epochs. The findings of this analysis can be observed in Table 5. When using DGL (Belilovsky, Eickenberg, and Oyallon 2019) as the baseline, after incorporating MLAAN, it only takes 300 epochs to reduce the Test Error by 45% for a Backbone of ResNet-32 (K=16) (He et al. 2016) and 52% for ResNet-110 (K=55) (He et al. 2016).
@@
Comparison of Features in Different Methods. To showcase the advanced capabilities of MLAAN, we conduct feature map analyses on different configurations, including DGL (Belilovsky, Eickenberg, and Oyallon 2019), DGL with MLM, DGL with LAM, and DGL with MLAAN. The resulting figures detailing these feature maps can be found in Figure 4.
@@
Decoupled Layer Accuracy Analysis. We have demonstrated that MLAAN achieves accuracy comparable to BP (Rumelhart et al. 1985) across different architectures. To further analyze MLAAN’s impact, we freeze the main network parameters and train a linear classifier for each gradient-isolated layer. Results are shown in Figure 5 using DGL (Belilovsky, Eickenberg, and Oyallon 2019) as the baseline.
@@
Representation Similarity Analysis. To further demonstrate our method’s effectiveness, we conduct Centered Kernel Alignment (CKA) (Kornblith et al. 2019) experiments. We incorporate MLAAN and calculate the CKA (Kornblith et al. 2019) similarity for each layer compared to BP (Rumelhart et al. 1985), then compute the mean values. Results are shown in Figure 6.
MLAAN significantly enhances DGL’s CKA similarity, particularly in early layers, aligning patterns more closely with BP (Rumelhart et al. 1985). The original method’s poor performance arises from early layers focusing on local optimization, which reduces overall performance.
@@ Conclusion
This paper introduces the Multilaminar Leap Augmented Auxiliary Network (MLAAN), a novel approach to resolving issues related to the lack of global features and myopic perspectives that lead to poor performance in existing supervised local learning. MLAAN consists of Multilaminar Local Modules (MLM) and Leap Augmented Modules (LAM). Our MLM addresses the issue of insufficient global features in existing work. We further propose that LAM enhance the auxiliary network for inter-module information exchange, effectively utilizing deeper information to mitigate shortsightedness. We integrate our MLAAN into two SOTA supervised local learning methods and evaluate the performance of different deep network architectures across four widely used datasets. The results demonstrate that our approach significantly enhances the performance of original supervised local learning methods, even surpassing that of E2E training.

### v6

@@ Introduction
End-to-end (E2E) backpropagation, a mainstream method for training deep neural networks, has achieved remarkable success in image segmentation (Wang, Peng, and Liu 2018), target tracking (Bao et al. 2022), visual sound separation (He et al. 2021), and other fields. Despite these accomplishments, E2E learning has its drawbacks. It necessitates retaining the complete computational graph and all intermediate activations post each local module’s computation (Wang et al. 2021a), increasing GPU memory usage. Additionally, the E2E structure delays parameter updates in initial layers until the entire forward and backward propagation cycles conclude, causing inefficiency (Jaderberg et al. 2017a), (Salimans et al. 2017). This sequential gradient processing also impedes model parallelization (Löwe, O’Connor, and Veeling 2019a), (Belilovsky, Eickenberg, and Oyallon 2020) and raises biological plausibility concerns due to its reliance on cross-layer backpropagation (Crick 1989), (Shen, Zhao, and Zeng 2022).
Due to the inherent flaws of E2E training (Illing et al. 2021), (Xiong, Ren, and Urtasun 2020a), some researchers are exploring local learning as an alternative. Local learning divides the network into gradient-isolated local modules and designs auxiliary networks to assist their training (Bengio et al. 2006). This method eliminates the need for sharing gradients and activation parameters between modules, reducing GPU memory requirements (Wang et al. 2021a). In addition, local learning uses local error signals, which align more closely with biological feasibility principles. However, the existing focus on local features and the lack of information exchange between local modules significantly lower overall performance than that of E2E training (Mostafa, Ramesh, and Cauwenberghs 2018). Despite some recent improvements in the performance of local learning (Belilovsky, Eickenberg, and Oyallon 2020), (Nøkland and Eidnes 2019), current methods still have myopia issues and insufficient global feature representation, limiting their application in precision-critical tasks. Therefore, the crucial goal of local learning is to ensure that each local module receives ample global information and facilitates information exchange between modules. Maintaining GPU memory efficiency while improving global performance is essential for making local learning a viable alternative to E2E training.
@@
Local learning is considered a promising alternative to traditional E2E methods (Nøkland and Eidnes 2019; Ma et al. 2024). However, the lack of local inter-module interaction and poor global performance has hindered its development (Bengio et al. 2006; Zhu et al. 2024). To improve global performance, some scholars have used self-supervised loss to preserve features during local module propagation. For example, Y. Xiong et al. (Xiong, Ren, and Urtasun 2020b) designed SimCLR loss for Local Contrastive Learning. Sindy et al. (Löwe, O’Connor, and Veeling 2019b) applied Comparative Predictive Coding loss to reduce feature loss in Greedy InfoMax. These methods perform well with a few network blocks but struggle with many blocks. Adrien et al.’s method (Journé et al. 2022), combining Hebbian learning with local learning, addresses this issue but faces challenges on large datasets. Methods like InfoPro (Wang et al. 2021a) and DGL (Belilovsky, Eickenberg, and Oyallon 2020) have improved local learning performance on large datasets such as ImageNet (Deng et al. 2009). However, due to gradient isolation, their performance still falls short of E2E methods (Su et al. 2024a, b). Enhancing inter-module communication to match or surpass E2E performance remains a critical issue.
@@
BP(ResNet-32)=19.35±0.24, BP(ResNet-110)=19.67±0.34
SVHN DGL 4.83±0.11 5.05±0.08 5.12±0.16 5.36±0.19
@@
Results on Image Classification Benchmarks.
We evaluate MLAAN’s performance, with results in Table 1. On CIFAR-10 (Krizhevsky, Hinton et al. 2009), increasing local modules enhances MLAAN’s benefits. For ResNet-32 (k=16) (He et al. 2016), the Test Error decreases from 14.08 and 12.93 to 6.36 and 6.37, marking a 55% and 51% improvement over the BP (Rumelhart et al. 1985) baseline. For ResNet-110 (k=55) (He et al. 2016), MLAAN with DGL (Belilovsky, Eickenberg, and Oyallon 2019) and InfoPro (Wang et al. 2021b) improves performance by 61% and 56%, respectively. On STL-10 (Coates, Ng, and Lee 2011), MLAAN achieves over 30% and 34% improvements under the same situations, while on SVHN (Netzer et al. 2011), gains exceed 42% and 49%.
@@
These results demonstrate MLAAN’s ability to improve deep learning models and outperform BP even with more local modules, reinforcing its effectiveness in improving accuracy and robustness in complex and deep architectures.
Memory Consumption.
We investigate the GPU memory usage of BP (Rumelhart et al. 1985) and MLAAN on CIFAR-10 (Krizhevsky, Hinton et al. 2009) and ImageNet (Deng et al. 2009), with results in Table 3.
For ResNet-32 and ResNet-110 (He et al. 2016), usage decreases by 22.3% and 70.2%, compared to BP (Rumelhart et al. 1985). ResNet-34, ResNet-101, and ResNet-152 (He et al. 2016) also show reductions of 15.4%, 13.1%, and 12.5%, respectively. These findings demonstrate that MLAAN performs better with less GPU memory usage than BP (Rumelhart et al. 1985).
@@
Comparison of Different Module Combinations.
We conduct experiments with ResNet-32 (K=16) (He et al. 2016) as the backbone, using DGL (Belilovsky, Eickenberg, and Oyallon 2019) as the baseline, to evaluate MLM, LAM, and MLAAN. The results are summarized in Table 4.
@@
Comparative Analysis of Unfair Epochs.
Undeniably, the utilization of MLAAN introduces additional Wall Time, prompting us to conduct a comparative analysis of unfair epochs. The findings can be observed in Table 5. When using DGL (Belilovsky, Eickenberg, and Oyallon 2019) as the baseline, after incorporating MLAAN, it only takes 300 epochs to reduce the Test Error by 45% for a Backbone of ResNet-32 (K=16) (He et al. 2016) and 52% for ResNet-110 (K=55) (He et al. 2016).
@@
Comparison of Features in Different Methods.
To showcase the advanced capabilities of MLAAN, we conduct feature map analyses on different configurations, including DGL (Belilovsky, Eickenberg, and Oyallon 2019), DGL with MLM, DGL with LAM, and DGL with MLAAN. The resulting figures can be found in Figure 4.
@@
Decoupled Layer Accuracy Analysis.
To further analyze MLAAN’s impact, we freeze the main network parameters and train a linear classifier for each gradient-isolated layer. Results are shown in Figure 5 using DGL (Belilovsky, Eickenberg, and Oyallon 2019) as the baseline.
@@
Representation Similarity Analysis.
We conduct Centered Kernel Alignment (CKA) (Kornblith et al. 2019) experiments. We incorporate MLAAN and calculate the CKA (Kornblith et al. 2019) similarity for each layer compared to BP (Rumelhart et al. 1985), then compute the mean values. Results are shown in Figure 6.
MLAAN significantly enhances DGL’s CKA similarity, particularly in early layers. The original method’s poor performance arises from early layers focusing on local optimization, which reduces overall performance.
@@ Conclusion
This paper introduces the Multilaminar Leap Augmented Auxiliary Network (MLAAN), a novel approach to resolving issues related to the lack of global features and myopic perspectives that lead to poor performance in existing supervised local learning. MLAAN consists of Multilaminar Local Modules (MLM) and Leap Augmented Modules (LAM). MLM addresses the issue of insufficient global features in existing work. We further propose that LAM enhance the auxiliary network for inter-module information exchange, effectively utilizing deeper information to mitigate shortsightedness. We integrate our MLAAN into two SOTA supervised local learning methods and evaluate the performance of various deep network architectures across four widely used datasets. The results demonstrate that our approach significantly enhances the performance of original supervised local learning methods, even surpassing that of E2E training.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (52278154) and the Natural Science Foundation of Jiangsu (BK20231429).
Appendix A Supplementary Material
Implement Detail
Our experiments use ResNet-32 and ResNet-110 (He et al. 2016) as backbone networks. We employ the SGD optimizer (Keskar and Socher 2017) with Nesterov momentum (Dozat 2016) at 0.8 and an L2 weight decay of 1e-4. ResNet-32 (He et al. 2016) is divided into 16 modules, while ResNet-110 (He et al. 2016) is segmented into 55 modules, each with its own parameters and auxiliary network.
Different hyperparameters are used for each dataset: batch size is 1024 for CIFAR-10 (Krizhevsky, Hinton et al. 2009) and SVHN (Netzer et al. 2011), and 128 for STL-10 (Coates, Ng, and Lee 2011). Training lasts 400 epochs with initial learning rates of 0.8 for CIFAR-10 (Krizhevsky, Hinton et al. 2009) and SVHN (Netzer et al. 2011), and 0.1 for STL-10 (Coates, Ng, and Lee 2011), following a cosine annealing scheduler (Loshchilov and Hutter 2016). For ImageNet (Deng et al. 2009), we use ResNet-34, ResNet-101, and ResNet-152 (He et al. 2016), training for 150 epochs with an initial learning rate of 0.05 and a batch size of 128. We maintain consistency with other training configurations as previously described for CIFAR-10 (Krizhevsky, Hinton et al. 2009).
Generalization Study
In this section, we aim to investigate the generalization performance of our proposed MLAAN method. To evaluate its effectiveness, we utilize the checkpoints trained on the CIFAR-10 (Krizhevsky, Hinton et al. 2009) and test them on the STL-10 (Coates, Ng, and Lee 2011), taking inspiration from previous work (Qu et al. 2021).
As shown in Table 6, we observe a notable disparity in the generalization abilities of DGL (Belilovsky, Eickenberg, and Oyallon 2019) and BP (Rumelhart et al. 1985). However, with the incorporation of our MLAAN method, we witness a significant improvement in test accuracy, surpassing even that of BP (Rumelhart et al. 1985). These results indicate that MLAAN, by facilitating information interaction among local modules, effectively enhances the generalization ability of supervised local learning methods.
Method ResNet-32(K=16) ResNet-110(K=55)
BP 35.98 36.78
DGL 31.95 33.16
DGL* 37.65 40.26
Table 6: Generalization study. Checkpoints are trained on the CIFAR-10 dataset and tested on the STL-10 dataset. The data in the table represents the test accuracy.
These findings emphasize the efficacy of our MLAAN method in improving the generalization capabilities of supervised local learning, ultimately leading to enhanced overall performance in the image classification task.